OpenStudy (anonymous):

I need some more exponential and logarithmic help ): Will list my two questions in the comments below.

OpenStudy (anonymous):

"Solve each question. Show all your work. Round your answer to four decimal places." A.)\[7^{4x}=415\]B.)\[\log _{7}(x+1)+\log _{7}(x-5)=1\]

OpenStudy (anonymous):

I'm not quite sure how to go about solving this, if someone could just refresh my memory and point me in the right direction on HOW to solve this, that would be fine.

OpenStudy (anonymous):